# New Numeric Invariants of an Unfolding of a Polycycle

### v1

@@ Abstract
In this paper new numeric invariants of structurally unstable vector fields in the plane are found. One of the main tools is an improved asymptotics of sparkling saddle connections that occur when a separatrix loop of a hyperbolic saddle breaks. Another main tool is a new topological invariant of two arithmetic progressions, both perturbed and unperturbed, on the real line. For the pairs of the unperturbed arithmetic progressions we give a complete topological classification.
@@ 1.2 New invariants
Let $v_{0}$ and $A$ be the same as before. Let $\gamma_{2}$ be the whole polycycle “tears of the heart”, and $\gamma_{1}\subset\gamma_{2}$ be the separatrix loop. Both polycycles are monodromic, $\gamma_{1}$ from the inside and $\gamma_{2}$ from the outside. Let $\Gamma$ be a transversal to the saddle loop, and $\Gamma_{1}$, $\Gamma_{2}$ be its interior and exterior parts with respect to $\gamma_{1}$. Let $\Delta_{1}\colon\Gamma_{1}\to\Gamma_{1}$ ($\Delta_{2}\colon\Gamma_{2}\to\Gamma_{2}$) be the monodromy map (the inverse monodromy map) corresponding to $\gamma_{1}$ (resp., $\gamma_{2}$). There exist $C^{1}$-smooth charts $x_{1,2}$ on $\Gamma_{1,2}$ such that the monodromy maps $\Delta_{j}$ have the form $x_{j}\mapsto C_{j}x_{j}^{\nu_{j}}$ in the coordinates $x_{1},x_{2}$ respectively, with $\nu_{1}=\lambda$ and $\nu_{2}=(\lambda^{2}\mu)^{-1}$. For a map $\Delta:x\mapsto Cx^{\nu}$ and $B>0$ we set
@@
$\Xi=\Bigl(\frac{\ln C_{1}}{1-\nu_{1}}-\ln B_{1}\Bigr)^{-1}\Bigl(\frac{\ln C_{2}}{1-\nu_{2}}-\frac{\ln C_{1}}{1-\nu_{1}}\Bigr).$
@@
Now consider two topologically equivalent families with ratios $\tau,\tilde{\tau}$ from Theorem 1. Since
$\tau-\tilde{\tau}=Ap+s,\quad s,p\in\mathbb{Z},$
one can prove (see Section 4.3 below) that
$\ln|\Xi|-\ln|\tilde{\Xi}|=s\ln\nu_{2},$
$\ln|\Theta|-\ln|\tilde{\Theta}|=p\ln\nu_{1}=-Ap\cdot\ln\nu_{2}.$
Thus, $\ln|\Xi|\pmod{\ln\nu_{2}}$ and $\ln|\Theta|\pmod{\ln\nu_{1}}$ describe the same new invariant.
@@ 1.4 Estimates of the monodromy mapping
Consider a local one-parametric unfolding of a monodromic hyperbolic polycycle such that only one separatrix connection of the original polycycle is broken with non-zero speed with respect to the parameter $\varepsilon$. Let $\nu<1$ be the product of the characteristic numbers of the saddles encountered along the polycycle. Then there exists a $C^{1}$-smooth chart for the first return map $f_{\varepsilon}$ of the perturbed polycycle such that in this chart we have
@@ 1.5 Definitions
Two families are strongly topologically equivalent if there exists a homeomorphism $H$ with above properties. Strong topological classification of families is quite rigid, whereas weak topological equivalence is too lousy, see e.g. the discussion in [3]. For this reason, we use here the notion of moderate topological equivalence. Whenever we write about equivalent families, we mean moderate topological equivalence.
@@ 2.2 Normalization of the monodromy map
Taking $k=N$ and $c=\min(1,s_{N}(0)/2)$, we get the required statement. Recall that we were working with $Q$ written in the normalizing chart of $\Delta_{1}$. We omit the straightforward check that, since $c<\Lambda(\varepsilon)$ and $c<1$, the map $Q$ will have the same form $Q(x,\varepsilon)=C(\varepsilon)x^{\Lambda(\varepsilon)}+O(x^{\Lambda(\varepsilon)+c})$ in any smooth $\varepsilon$-dependent coordinate, but with a different $C(\varepsilon)$.
@@ 3.2 Asymptotics of the sparkling saddle connections: the model case
In the case of a monodromic polycycle (with the product of characteristic numbers of saddles $\Lambda(\varepsilon)<1$) perturbed in a one-parameter family where only one connection is broken, the sequence $(\varepsilon_{n})$ of parameter values that correspond to separatrix connections with an external saddle has the form
@@ 3.5 Topology of two pairs of slightly perturbed arithmetic progressions
$As+p=As_{2}+p_{2}$
so $(s-s_{2})A=p_{2}-p$, which would imply $A$ rational, contradicting the irrationality of $A$.
@@ 4.1 Sketch of the proof
Let $((i_{n}),(e_{n}))$ and $((\tilde{i}_{n}),(\tilde{e}_{n}))$ be the pairs of exponentially perturbed arithmetic progressions generated by two shift-exp-Liouvillian families. If these pairs are topologically equivalent at infinity, then the bases of the sequences $(i_{n})$ and $(\tilde{i}_{n})$ coincide, the bases of $(e_{n})$ and $(\tilde{e}_{n})$ coincide, and the logarithms of their relative scale coefficients coincide modulo the logarithm of the base of $(e_{n})$ and $(\tilde{e}_{n})$.
@@
Let two shift-exp-Liouvillian families be moderately topologically equivalent. Then the pairs of sequences $((i_{n}),(e_{n}))$ and $((\tilde{i}_{n}),(\tilde{e}_{n}))$ generated by these families are topologically equivalent. Then Lemma 4.1 yields $\lambda=\tilde{\lambda}$. Since the parameter $A=\frac{-\ln\lambda}{\ln(\lambda^{2}\mu)}$ must be the same for the equivalent families, the equality $\mu=\tilde{\mu}$ follows. This completes the proof of Theorem 2. Likewise, by Lemma 4.1 the logarithms of the relative scale coefficients coincide modulo $\ln(\lambda^{2}\mu)$, which yields Theorem 3. ∎
@@ 4.2 Shift-exp-Liouvillian three-parameter families
$\Xi=\Bigl(\frac{\ln C_{1}}{1-\nu_{1}}-\ln B_{1}\Bigr)^{-1}\Bigl(\frac{\ln C_{2}}{1-\nu_{2}}-\frac{\ln C_{1}}{1-\nu_{1}}\Bigr).$
@@ 4.4 Shift-exp-Liouvillian families are generic
Define the subsets of $\mathbb{R}$.
@@
Openness is immediate: small perturbations of a family change its parameters only slightly, so an inclusion

### v2

# New Numerical Invariants of an Unfolding of a Polycycle “Tears of the Heart”

@@ Abstract
In this paper new numerical invariants of structurally unstable vector fields in the plane are found. One of the main tools is an improved asymptotics of sparkling saddle connections that occur when a separatrix loop of a hyperbolic saddle breaks. Another main tool is a new topological invariant of two arithmetic progressions, both perturbed and unperturbed, on the real line. For the pairs of the unperturbed arithmetic progressions we give a complete topological classification.
@@ 1.2 New invariants
Let $v_{0}$ and $A$ be the same as before. Let $\gamma_{2}$ be the whole polycycle “tears of the heart”, and $\gamma_{1}\subset\gamma_{2}$ be the separatrix loop. Both polycycles are monodromic, $\gamma_{1}$ from the inside and $\gamma_{2}$ from the outside. Let $\Gamma$ be a transversal to the saddle loop, and $\Gamma_{1}$, $\Gamma_{2}$ be its interior and exterior parts with respect to $\gamma_{1}$. Let $\Delta_{1}\colon\Gamma_{1}\to\Gamma_{1}$ ($\Delta_{2}\colon\Gamma_{2}\to\Gamma_{2}$) be the monodromy map (the inverse monodromy map) corresponding to $\gamma_{1}$ (resp., $\gamma_{2}$). There exist $C^{1}$-smooth charts $x_{1,2}$ on $\Gamma_{1,2}$ such that the monodromy maps $\Delta_{j}$ have the form $x_{j}\mapsto C_{j}x_{j}^{\nu_{j}}$ in the coordinates $x_{1},x_{2}$ respectively, with $\nu_{1}=\lambda$ and $\nu_{2}=(\lambda^{2}\mu)^{-1}$. For a map $\Delta:x\mapsto Cx^{\nu}$ and a small $B>0$ we set
@@
$\Xi=\Bigl(\frac{\ln C_{1}}{1-\nu_{1}}-\ln B_{1}\Bigr)^{-1}\Bigl(\frac{\ln C_{2}}{1-\nu_{2}}-\frac{\ln C_{1}}{1-\nu_{1}}\Bigr)$
@@
Now consider two topologically equivalent families with ratios $\tau,\tilde{\tau}$ in Theorem 1 such that
$\tau-\tilde{\tau}=As+p,\quad s,p\in\mathbb{Z}.$
One can prove (see Sec. 4.3 below) that
$\ln|\Xi|-\ln|\tilde{\Xi}|=s\ln\nu_{1}.$
Then a direct calculation shows that
$\ln|\Theta|-\ln|\tilde{\Theta}|=p\ln\nu_{2}.$
Thus, $\ln|\Xi|\pmod{\ln\nu_{1}}$ and $\ln|\Theta|\pmod{\ln\nu_{2}}$ are representations of the same new invariant.
@@ 1.4 Estimates of the monodromy mapping
Consider a local one-parametric unfolding of a monodromic hyperbolic polycycle such that only one separatrix connection of the original polycycle is broken with non-zero speed with respect to the parameter $\varepsilon$. Let $\nu<1$ be the product of the characteristic numbers of the saddles encountered along the polycycle and assume that these characteristic numbers are irrational. Then there exists a $C^{1}$-smooth chart for the first return map $f_{\varepsilon}$ of the perturbed polycycle such that in this chart we have
@@ 1.5 Definitions
Two families are strongly topologically equivalent if there exists a homeomorphism $H$ with above properties. Strong topological classification of families is quite rigid, whereas weak topological equivalence is too lousy, see e.g. the discussion in [3]. The results of [1] that we use are stated using the notion of moderate topological equivalence. Hence, whenever we write about equivalent families, we mean the moderate topological equivalence. However, we believe that the moderate equivalence can be replaced by the weak equivalence in all the results.
@@ 2.2 Normalization of the monodromy map
Taking $k=N$ and $c<\min(1,s_{N}(0)/2)$, we get the required statement. Recall that we were working with $Q$ written in the normalizing chart of $\Delta_{1}$. We omit the straightforward check that, since $c<\Lambda(\varepsilon)$ and $c<1$, the map $Q$ will have the same form $Q(x,\varepsilon)=C(\varepsilon)x^{\Lambda(\varepsilon)}+O(x^{\Lambda(\varepsilon)+c})$ in any smooth $\varepsilon$-dependent coordinate, but with a different $C(\varepsilon)$.
@@ 3.2 Asymptotics of the sparkling saddle connections: the model case
In the case of a monodromic polycycle (with irrational characteristic numbers that have product $\Lambda<1$) perturbed in a one-parameter family where only one connection is broken, the sequence $(\varepsilon_{n})$ of parameter values that correspond to separatrix connections with an external saddle has the form
@@ 3.5 Topology of two pairs of slightly perturbed arithmetic progressions
$As-p=As_{2}-p_{2}$
so $(s-s_{2})A=p-p_{2}$, which would imply $A$ rational, contradicting the irrationality of $A$.
@@ 4.1 Sketch of the proof
Let $((i_{n}),(e_{n}))$ and $((\tilde{i}_{n}),(\tilde{e}_{n}))$ be the pairs of exponentially perturbed arithmetic progressions generated by two shift-exp-Liouvillian families. If these pairs are topologically equivalent at infinity, then the bases of the sequences $(i_{n})$ and $(\tilde{i}_{n})$ coincide, the bases of $(e_{n})$ and $(\tilde{e}_{n})$ coincide, and the logarithms of their relative scale coefficients coincide modulo the logarithm of the base of $(i_{n})$ and $(\tilde{i}_{n})$.
@@
Let two shift-exp-Liouvillian families be moderately topologically equivalent. Then the pairs of sequences $((i_{n}),(e_{n}))$ and $((\tilde{i}_{n}),(\tilde{e}_{n}))$ generated by these families are topologically equivalent. Then Lemma 4.1 yields $\lambda=\tilde{\lambda}$. Since the parameter $A=\frac{-\ln\lambda}{\ln(\lambda^{2}\mu)}$ must be the same for the equivalent families, the equality $\mu=\tilde{\mu}$ follows. This completes the proof of Theorem 2. Likewise, by Lemma 4.1 the logarithms of the relative scale coefficients coincide modulo $\ln\lambda$, which yields Theorem 3. ∎
@@ 4.2 Shift-exp-Liouvillian three-parameter families
$\Xi=\Bigl(\frac{\ln C_{1}}{1-\nu_{1}}-\ln B_{1}\Bigr)^{-1}\Bigl(\frac{\ln C_{2}}{1-\nu_{2}}-\frac{\ln C_{1}}{1-\nu_{1}}\Bigr)$
@@ 4.4 Shift-exp-Liouvillian families are generic
Define the subsets of $Q^{+}_{N,q}(\gamma,u,\Xi,\lambda)\subset\mathbb{R}$ by
@@
Openness is immediate: mall perturbations of a family change its parameters only slightly, [^10] Note that the canonical chart depends continuously on the family of vector fields, see Theorem A in P. Berger and B. Reinke, arXiv:2210.05256, 2022. so an inclusion